OpenStudy (redrose5004):

A class of 12 students has taken an exam, and the mean of their scores is 71. One student takes the exam late, and scores 92. After including the new score, what is the mean score for all 13 exams? If you get a decimal answer, you should round to the nearest tenth.

OpenStudy (kropot72):

Total score for 12 students = 12 * 71 Total score for 13 students = (12 * 71) + 92 = 944 Mean score for 13 students = 944/13 = 72.6